OpenStudy (anonymous):

Barry deposited $4000 into two saving accounts bearing simple interest. One of the accounts has an interest rate of 10% while the other rate is 4%. If the total interest earned after one year is $352, find the amount deposited into each of the accounts.

OpenStudy (anonymous):

let his savings in the accounts be x and 4000-x then the interest on 1 yr from the are respectively 10x /100 and 4*(4000-x)/100 hence by the problem we have x/10 +(16000-4x)/100=352 or x/10 +160 -4x/100 =352 or (10x-4x)/100 =352-160 or 6x/100=192 or x=192/6 *100=3200 thus he invested 3200 in 10% account and (4000-3200)=800 in 4%